O Determine he arnat Cliaegh of all he atrms except hydroen 6, which of i r frules are kalp organic compourd: ppose a shucture for a nolecula that only had for Carin - atome as nedad) ④ Are thefollowing moleculen, resonance frmrof each other reamance torm.s and
media%2F470%2F47000836-2e07-4dcb-aca1-03
media%2F2b1%2F2b1230c1-8c0d-4406-9f4c-f3
Answer #1

1. Outline: The formal charge on an atom = NA - NM , where NA = no. of electrons in the valence shell of free atom and NM = no. of electrons belonging to the atom in the molecule, NM = NLP - (1/2)NBP ; NLP and NBP are the no. of electrons in the lone pairs and bonding pairs, respectively.

a) The formal charge on N in NH3-BH3 = 5 - (0 + (1/2)*8) = +1

The formal charge on B in NH3-BH3 = 3 - (0 + (1/2)*8) = -1

b) The formal charge on all the C's = 4 - (0 + (1/2)*8) = 0

The formal charge on N = 5 - (0 + (1/2)*8) = +1

The formal charge on O = 6- (6 + (1/2)*2) = -1